Use the Hardy-Weinberg equations (p2 + 2pq + q2 = 1 and p + q = 1) to find the genotypes of the following hypothetical phenotypic information:

a.       Within a particular population, you observe that 450 individuals have freckles (dominant, F) and 550 do not have freckles (recessive, f).

b.      Homozygous dominant (frequency) ______(A)_________

c.       Heterozygous (frequency) ______(B)_____________

d.      Homozygous recessive (frequency) _______(C)____________
Answer #1

Answer:

Frequency of freckles = q^2= 550/1000= 0.55

Frequency of recessive allele =q= sqrt of 0.55= 0.74

Frequency of dominant allele= p=1- 0.74= 0.26

Frequency of Homozygous dominant= pp= 0.26*0.26= 0.07

Frequency of heterozygote =pq= 2*0.26*0.74= 0.38

Frequency of Homozygous recessive= 0.74*0.74= 0.55